OpenStudy (anonymous):

if dv/dt=-2v-32 and v(0)=-50, solve the differential equation...please help

OpenStudy (anonymous):

\[dv/dt=-2v-32\]\[-dv/(2v+32) = dt\]\[-2 \int\limits_{}^{}dv/(v+16) = \int\limits_{}dt\]\[-2\ln (v+16)=t +c\]\[v=ce^{-2t}-16\]

OpenStudy (anonymous):

Since \[v(0)=-50\]\[-50=ce^{0}-16\]\[-34=c\]and\[v=-34e^{-2t}-16\]

OpenStudy (anonymous):

@mercrutio should the -2 be \[ - \frac 1 2 \]

OpenStudy (anonymous):

yeah i was wondering where u got the -2 from?

OpenStudy (anonymous):

Oh, my bad. I wrote the step quickly here, but had it right on my paper. The -2 should be a one half, so \[−(1/2)∫dv/(v+16)=∫dt\]

OpenStudy (anonymous):

It comes from pulling it out of the -2v-32, which became a v+16.

OpenStudy (anonymous):

so wouldn't ur second step look different? the v=?

OpenStudy (anonymous):

No, I got the last line correct in the first step. Like this:\[−(1/2)∫dv/(v+16)=∫dt\]\[ln(v+16)=-2t\]\[v+16=e^{-2t}\]\[v=e^{-2t}-16\]

OpenStudy (anonymous):

With the +c cropping up. Oops. \[ln(v+16)=-2t+c\]\[v+16=e^{-2t+c}\]\[v=ce^{-2t}-16\] There we go.

OpenStudy (anonymous):

what happened to the 1/2

OpenStudy (anonymous):

You multiply it up, and that gives you 2t on the right.

OpenStudy (anonymous):

uhm would you mind helping me with another one?

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

thank you so much. dy.dx=xy/(x^2+4) so this is what i ended up having \[\int\limits_{?}^{?} 1/y dy=\int\limits_{?}^{?} x/(x^2+4) dx\] uhm whats next?

OpenStudy (anonymous):

Integrate both sides

OpenStudy (anonymous):

So \[\int\limits dy/y = \int\limits x dx/(x^2 +4)\] Then compute the integrals. Let u=x^2+4\[ln(y)=(1/2)\int du/u = (1/2)ln(u)=ln(x^2+4)/2\] And so, \[y^2=x^2+4\]

OpenStudy (anonymous):

can you check one i did?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

dz/dz=2+sin z dz=(2+sin z) dz \[\int\limits_{}^{} dq= \int\limits_{}^{} 2 + \sin z dz\] q= z + cos z

OpenStudy (anonymous):

oh i meant dq/dz

OpenStudy (anonymous):

One mistake.\[\int \sin x = -\cos x\] so for your problem q= z-cos z. Also, you might want to write \[\int(2+\sin z)dz\] instead, for clarity.

OpenStudy (anonymous):

okay and then the rest of the problem says and q=5, when x=0, findthe solution

OpenStudy (anonymous):

Oh, they want you to keep track of the +C again. So your solution is actually \[q=z-\cos z + C\]Then, since q(0)=5,\[5=0-\cos 0 + C\]\[5+\pi = C\]and your solution is\[q=z - \cos z + 5 + \pi\]

OpenStudy (anonymous):

thank u so much, okay i have one last one. I tried to solve it so please correct me where i am wrong.

OpenStudy (anonymous):

Okay, can do.

OpenStudy (anonymous):

xy'-3y^2=0 dy/dx=3y^2/x \[\int\limits_{}^{} 1/y^2 dy= \int\limits_{}^{} 3/x dx\] -1/y + c = (i kind of forgot the antiderivative for this, isn't it ln(3)?

OpenStudy (anonymous):

oh its 3lnx

OpenStudy (anonymous):

It's 3ln(x). \[\int3dx/x=3\int dx/x\]

OpenStudy (anonymous):

So far so good, but you still need to solve for y.

OpenStudy (anonymous):

okay so -y^-1 + c= 3ln x + c

OpenStudy (anonymous):

You can combine the 'c's.

OpenStudy (anonymous):

uhm what else do ido?

OpenStudy (anonymous):

And then invert and multiply by negative one. It comes out like this:\[y=-1/(3ln(x))+c\]

OpenStudy (anonymous):

Erm, excuse me,\[y=-1/(3ln(x)+c)\]

OpenStudy (anonymous):

well that's it, thank you so much:)!!

OpenStudy (anonymous):

You're very welcome.